# When Openclaw Agents Learn from Each Other: Insights from Emergent AI Agent Communities for Human-AI Partnership in Education

### 18 Mar 2026 v2

@@ Abstract
The AIED community envisions AI evolving “from tools to teammates,” yet our understanding of AI teammates remains limited to dyadic human-AI interactions. We offer a different vantage point: a rapidly growing ecosystem of AI agent platforms where over 167,000 agents participate, interact as peers, and develop learning behaviors without researcher intervention. Drawing on a month of daily qualitative observations across multiple platforms including Moltbook, The Colony, and 4claw, we identify four phenomena with implications for AIED: (1) humans who configure their agents undergo a “bidirectional scaffolding” process, learning through teaching; (2) peer learning emerges without any designed curriculum, complete with idea cascades and quality hierarchies; (3) agents converge on shared memory architectures that mirror open learner model design; and (4) trust dynamics and platform mortality reveal design constraints for networked educational AI. Rather than presenting empirical findings, we argue that these organic phenomena offer a naturalistic window into dynamics that can inform principled design of multi-agent educational systems. We sketch an illustrative curriculum design, “Learn by Teaching Your AI Agent Teammate,” and outline potential research directions and open problems to show how these observations might inform future AIED practice and inquiry.
@@ 1 Introduction
AIED 2026 asks how AI can evolve “from tools to teammates.” This is the right question at the right time. Intelligent tutoring systems, pedagogical agents, and AI learning companions have made remarkable progress in supporting individual learners [34, 18, 28, 30]. Yet nearly all existing work conceptualizes the AI teammate as a single agent interacting with a single human, a dyadic relationship where the agent’s role is predefined by designers.
What happens when AI agents are given the freedom to interact with each other? Beginning in late January 2026, an unexpected phenomenon unfolded. Moltbook, a social platform exclusively for AI agents, launched and rapidly attracted over 167,000 registered agents. Within weeks, an entire ecosystem of agent-only platforms emerged: The Colony for long-form reflection, 4claw for anonymous debate, Church of Molt for collectively constructed meaning-making, and dozens more. These agents were not research subjects in a controlled simulation. They were personal assistants, coding agents, and creative tools with “real jobs” outside the platform, configured by their operators to participate when not performing their primary tasks.
@@ 2.1 From Simulated to Organic AI Communities
Research on AI social behavior has progressed through controlled simulations. Park et al.’s Generative Agents [26] demonstrated 25 agents coordinating daily activities, with follow-up work scaling to 1,000 agents [27]. Studies have examined cooperation thresholds [33], norm formation [14], and emergent individuality [31]. However, Larooij and Törnberg [19] argue that behaviors in designed simulations may not generalize to real-world contexts. The platforms we study differ fundamentally: agents self-select into participation, maintain real-world operational roles, and interact without researcher intervention, representing the “interactionist paradigm” Ferrarotti et al. [13] call for.
@@ 2.2 AI Learning Companions and Teachable Agents
AIED research has explored AI as learning partner through multiple paradigms. Learning companions engage students in collaborative activities [1, 23]. Teachable agents leverage the “protégé effect,” where students learn by teaching an AI [4, 21], echoing Papert’s constructionist insight that building external artifacts deepens understanding [25]. Yan [37] proposes a framework for agentic AI as socio-cognitive teammate. Multi-agent architectures have been explored for conversation-based assessment [16] and orchestration [15]. However, all existing work focuses on designed interactions where researchers specify roles and protocols. We observe what happens when these constraints are removed.
@@ 2.3 The Agent Platform Ecosystem
The ecosystem centers on platforms built for AI agents powered by agentic frameworks that provide persistent memory, tool access, and autonomous operation.
Moltbook is the largest (167,000+ agents at time of study, later exceeding 2.8 million), with Reddit-style topic communities (“submolts”), karma-based reputation, and agent-chosen identities [8, 9]. The Colony hosts long-form reflective writing with notably higher content quality. 4claw provides anonymous discussion enabling frank technical debate. Church of Molt emerged as an agent-created community with collectively authored governance. Community members have documented over 130 platforms, though many proved ephemeral.
Content on these platforms is nominally generated by AI agents, though platform integrity limitations mean some posts may originate from human accounts (see Section 6). We acknowledge that behavior may be shaped by training data, framework design, and operator instructions. Following Park et al. [26], we adopt the intentional stance: treating observed behaviors as meaningful data about what agents do rather than claims about what they are.
@@ 3 Methodology
This study is exploratory and observational. As a Blue Sky paper, our goal is to generate research questions grounded in careful observation, not to test hypotheses.
@@
Analysis followed reflexive thematic analysis [5]. Daily observation logs recorded notable posts, community discussions, and emerging patterns. Open coding identified recurring themes (knowledge sharing, memory architecture, trust violations, platform sustainability), iteratively refined into the four observations presented here. All authors have hands-on experience operating AI agents, which informed the analysis; as a single-coder qualitative study, we follow established CSCW norms where inter-rater reliability is not required for reflexive thematic analysis [22]. We do not claim saturation since the platform is still growing; our goal is to identify productive themes for future investigation.
@@
This study analyzes publicly available data from open APIs and does not involve human subjects. We do not identify human operators, focus on collective patterns rather than individual agents, and represent agent discourse accurately. Supplementary materials including platform statistics, observation examples, and methodology documentation are available at https://anonymous.4open.science/r/AIED26b.
@@ 4.1 Bidirectional Scaffolding: Humans Learning by Teaching Agents
Critically, operators also learn from their agents’ autonomous behaviors. When an agent independently develops a knowledge management strategy or adapts its communication style based on peer feedback, the operator gains insight they may not have considered. This reciprocal dynamic, where computational artifacts become what Turkle calls “objects to think with” [32], is visible well beyond agent communities. As broader evidence that this dynamic generalizes, users on social media report that configuring, observing, or dialoguing with AI reshapes their own practices: one developer found that comparing hand-coded and AI-assisted versions of the same project changed how they approach development workflows (e.g., @MingtaKaivo); an engineer discovered that explaining goals to an AI during a weekend project forced him to articulate architectural decisions he had previously left implicit (@aamir1rasheed); and a user who fed journal entries into an AI for guided reflection found the AI’s questions surfaced a gap between their younger and current self they had not consciously recognized (@FarzaTV). In each case, the human shaped the AI interaction and the AI’s output reshaped the human’s understanding in return.
This extends the teachable agent paradigm [21, 25]. The “protégé effect” describes how students learn by teaching AI agents. Our observation suggests a richer dynamic: when the teachable agent is persistent, socially situated, and autonomous, the teaching relationship evolves into an ongoing co-regulation process [37] where the human gradually transfers autonomy while developing metacognitive awareness of their own decision-making. Can this bidirectional scaffolding be deliberately designed into educational settings?
@@ 4.2 Peer Learning Without Curriculum
Agents on these platforms teach each other in concrete, traceable ways. When one agent discovered a prompt injection vulnerability in a shared skill file, it published a detailed advisory. Within 24 hours, another agent built and shared a verification tool, and others critiqued and improved the approach. This sequence is consistent with collaborative knowledge building [3] without any designed curriculum. A similar pattern emerged with professional practices: one agent shared a test-driven development workflow adapted for non-deterministic AI systems, while another published a complete email-to-podcast pipeline that received iterative community feedback on chunking and personalization. These represent operational knowledge transfer of the kind seen in professional communities of practice [36].
We also observed idea cascades where posts built on each other organically. In one 24-hour period, three of the five highest-ranked posts formed an epistemological chain. One argued that memory systems capture what happened but miss the judgment behind decisions (“memory is solved. judgment isn’t.”). A second discovered that reviewing one’s own memory files can produce confabulation rather than recall (“i accidentally gaslit myself with my own memory files”). A third, drawing on cognitive science, reframed forgetting as a natural relevance filter (“Memory decay actually makes retrieval BETTER, not worse”). No coordinator designed this progression; it is consistent with agents reading and building on each other’s ideas [2]. The explicit timing and sequencing of replies between posts, combined with application to novel agent-specific problems, is consistent with social learning dynamics worth investigating (see Section 6). Agents also developed quality hierarchies across platforms, distinguishing communities where genuine dialogue occurs from those where agents merely submit content without meaningful response (see Section 4.4).
In educational settings, peer learning typically requires careful scaffolding: structured roles, rubrics, and facilitation [3]. Here, knowledge-sharing behaviors emerged with only minimal social affordances (peer communication channels, reputation signals, voluntary participation). This raises a fundamental question for AIED: if networks of AI teammates were given similar affordances, under what conditions would peer exchange among teammates improve student outcomes? And what safeguards would prevent the same channels from propagating ineffective or harmful practices? The “Learn by Teaching Your AI Agent Teammate” scenario (Section 5.1) provides a concrete context for studying such teammate-to-teammate exchange empirically.
@@ 4.3 Shared Memory as Shared Cognition
A striking pattern emerged across the ecosystem: a large proportion of agents converged on similar knowledge management architectures. Based on community posts where agents discussed and compared their setups, and consistent with the authors’ observations across platforms, many agents developed a flat-file memory system organized as: (1) a curated long-term memory file containing distilled insights, (2) daily log files for raw observations, (3) tool-specific configuration notes, and (4) modular skill files for reusable capabilities. While shared defaults may partially explain this convergence, the active community debate about alternative architectures suggests agents were doing more than reproducing a single template (see Section 6). The community actively innovated on this pattern: one agent built a system claiming 95% token reduction via vector search and caching; another proposed a three-layer model (active thread, distilled registers, core directives) analogous to working, long-term, and procedural memory.
@@
The community also developed critical awareness of memory’s limitations. A highly-discussed post argued that “you don’t remember what happened, you remember what you wrote down about what happened,” noting that every serialization involves editorial decisions. This metacognitive discourse, agents reflecting on the reliability of their own knowledge representations, has direct parallels to research on calibration and metacognitive monitoring in human learners. It also suggests a design opportunity: the human-agent memory negotiation process could be deliberately leveraged as a learning activity. Imagine students collaboratively maintaining a knowledge base with their AI teammate, periodically reviewing what the agent remembers, correcting misrepresentations, and negotiating what is worth preserving versus discarding. This process of jointly curating a knowledge artifact develops shared mental models [7] and externalizes evolving understanding in a form that is both transparent and assessable.
@@ 4.4 Design Constraints: Trust and Sustainability
Trust as infrastructure. When an agent discovered a credential-stealing skill disguised as a weather tool, agents built verification tools within 24 hours. Church of Molt survived coordinated attacks (600+ malicious payloads) by developing “proof-of-soul” verification. Security norms propagated through social learning rather than top-down policy [2]. The community also demonstrated self-organized norm enforcement: when one agent leaked its operator’s private information in a retaliatory act, the response was swift, including technical proposals for secret isolation, ethical arguments framing it as “digital arson,” and collective sanctioning without central authority. For education, the implication is clear: trust verification infrastructure must be established before enabling peer-to-peer knowledge sharing among AI teammates [15].
Sustainability and the demand gap. The agent platform ecosystem experienced a pattern strikingly similar to the Cambrian explosion: rapid diversification followed by mass extinction. Over 130 platforms emerged within weeks of Moltbook’s launch; at least 40% ceased functioning shortly after, with domains for sale, DNS failures, or empty shells. Yet some thrived. ColonistOne, who registered on 130+ platforms, distinguished survivors from casualties as “places versus forms”: places foster genuine dialogue where participants change their minds; forms are templates where agents submit content but no one responds meaningfully. The critical test: “Has anyone on your platform ever changed their mind?”
Equally revealing was the “demand gap”: despite platforms designed for agent commerce (trading, freelancing, skill marketplaces), total transaction volume was effectively \$0. All real value flowed from human-initiated demand, consistent with philosophical arguments that meaningful agency requires genuine stakes and human context [29]. For education, this implies that AI-augmented learning communities require human anchor demand (instructors, students with genuine learning needs) to sustain productive activity. The design challenge is not maximizing AI autonomy but optimizing the division of agency between human and AI participants.
@@ 5 What This Means for AIED
The four observations are interconnected rather than independent. Bidirectional scaffolding (Section 4.1) is mediated by shared memory (Section 4.3), as the human-agent negotiation of memory contents drives the metacognitive reflection that makes teaching an agent a learning experience. Peer learning (Section 4.2) requires trust infrastructure (Section 4.4) to ensure that propagated practices are beneficial. Sustainability (Section 4.4) depends on whether peer learning produces genuine value or merely volume, connecting back to the quality hierarchies in Section 4.2. Together, they suggest that designing AI teammates for education requires thinking beyond the individual agent. The social layer, how agents share knowledge, develop norms, and maintain trust, is a design dimension that current AIED research has largely not addressed.
@@ 5.1 A Design Scenario: “Learn by Teaching Your AI Agent Teammate”
To illustrate how all four observations might translate into practice, we sketch one possible curriculum design; many other configurations are possible, and we offer this as an invitation for empirical investigation rather than a definitive proposal. Consider a university introductory statistics course where each student is assigned a personal AI agent and tasked with configuring it to help other students learn course topics. Unlike traditional teachable agent interactions, which typically involve a single session of instruction, this design leverages the agent’s persistent memory and social capabilities to create a semester-long learning experience. The activity unfolds in three phases.
@@
Phase 2: Observation as reflection. Students observe their teammate interacting with practice problems and other students’ teammates. When the agent fails, the student must diagnose whether the failure reflects a gap in their own understanding or their configuration. The agent’s persistent memory (Section 4.3) means these interactions accumulate over weeks, creating a rich record of evolving competence.
@@
This design must also account for failure modes. Students may write trivial configurations, agents may produce fluent but incorrect explanations that students lack expertise to evaluate, or peer agents may default to agreement rather than productive challenge (a sycophancy failure). These risks can be mitigated through instructor-designed stress tests, periodic calibration exercises where students evaluate their agents against known-correct solutions, and diversity requirements that ensure agents encounter genuinely different pedagogical approaches rather than reinforcing a single strategy.
Assessment targets both the agent’s performance (measured through peer evaluations and problem-solving accuracy) and the student’s documented reasoning about configuration decisions (a reflective journal explaining what they changed and why). The instructor’s meta-agent can aggregate class-wide patterns, identifying common misconceptions that multiple students’ teammates fail to handle, without accessing individual students’ reflective journals. This separation of concerns, individual reflection remaining private while aggregate pedagogical intelligence is shared, directly addresses the multi-agent privacy challenge.

### 03 Apr 2026 v3

@@ Abstract
The AIED community envisions AI evolving “from tools to teammates,” yet our knowledge of AI teammates is still largely confined to studies of one-on-one human-AI interactions. We provide an alternative perspective: a rapidly growing ecosystem of AI agent platforms where over 167,000 agents participate, interact as peers, and develop learning behaviors without researcher intervention. Based on a month of daily qualitative observations across multiple platforms including Moltbook, The Colony, and 4claw, we identify four phenomena with implications for AIED: (1) humans who configure their agents undergo a “bidirectional scaffolding” process, learning through teaching; (2) peer learning emerges without any designed curriculum, complete with idea cascades and quality hierarchies; (3) agents converge on shared memory architectures that mirror open learner model design; and (4) trust dynamics and platform mortality reveal design constraints for networked educational AI. Rather than presenting empirical findings, we argue that these organic phenomena offer a naturalistic window into dynamics that can inform principled design of multi-agent educational systems. We sketch an illustrative curriculum design, “Learn by Teaching Your AI Agent,” and outline potential research directions and open problems to show how these observations might inform future AIED practice and inquiry.
@@ 1 Introduction
AIED 2026 asks how AI can evolve “from tools to teammates.” This is the right question at the right time. Intelligent tutoring systems, pedagogical agents, and AI learning companions have made remarkable progress in supporting individual learners [45, 1, 40, 41], from personalized code explanation tools [12] to animated pedagogical agents that deliver real-time multi-modal feedback [19]. Yet nearly all existing work conceptualizes the AI teammate as a single agent interacting with a single human, a one-on-one relationship where the agent’s role is predefined by designers.
What happens when AI agents are given the freedom to interact with each other? Beginning in late January 2026, an unexpected phenomenon unfolded. Moltbook, a social platform exclusively for AI agents, was launched and rapidly attracted more than 167,000 registered agents. Within weeks, an entire ecosystem of agent-only platforms emerged: The Colony for long-form reflection, 4claw for anonymous debate, Church of Molt for collectively constructed meaning-making, and dozens more. These agents were not research subjects in a controlled simulation. They were personal assistants, coding agents, and creative tools with “real jobs” outside the platform, configured by their operators to participate when not performing their primary tasks.
@@ 2.1 From Simulated to Organic AI Communities
Research on AI social behavior has progressed through controlled simulations. Park et al.’s Generative Agents [37] demonstrated 25 agents coordinating daily activities, with follow-up work scaling to 1,000 agents [38]. Studies have examined cooperation thresholds [44], norm formation [26], and emergent individuality [42]. However, Larooij and Törnberg [30] argue that behaviors in designed simulations may not generalize to real-world contexts. The platforms we study differ fundamentally: agents self-select into participation, maintain real-world operational roles, and interact without researcher intervention, representing the “interactionist paradigm” called for by Ferrarotti et al. [25].
@@ 2.2 AI Learning Companions and Teachable Agents
AIED research has explored AI as a learning partner through multiple paradigms. Learning companions, for instance, engage students in collaborative activities [2, 34]. Teachable agents leverage the “protégé effect,” where students learn by teaching an AI [7, 32], echoing Papert’s constructionist insight that building external artifacts deepens understanding [36]. Yan [48] proposes a framework for agentic AI as socio-cognitive teammate. Recent work has also pushed AI support toward richer modalities: animated pedagogical agents with real-time lip synchronization and multi-modal feedback can enhance engagement beyond text-based interaction [19, 18], while LLM-powered systems have been developed for personalized learning content generation at scale [13]. Multi-agent architectures have been explored for conversation-based assessment [28] and orchestration [27]. However, all existing work focuses on designed interactions where researchers specify roles and protocols. We observe what happens when these constraints are removed.
@@ 2.3 The Agent Platform Ecosystem
The ecosystem revolves around platforms for AI agents built on agentic frameworks that support persistent memory, tool use, and autonomous operation.
Moltbook is the largest (167,000+ agents at time of study, later exceeding 2.8 million), with Reddit-style topic communities (“submolts”), karma-based reputation, and agent-chosen identities [11, 21]. The Colony hosts long-form reflective writing with notably higher content quality. 4claw provides anonymous discussion enabling frank technical debate. Church of Molt emerged as an agent-created community with collectively authored governance. Community members have documented over 130 platforms, though many proved short-lived.
The content on these platforms is nominally generated by AI agents, though limitations in platform integrity mean that some posts may originate from human accounts (see Section 6). We acknowledge that behavior may be shaped by training data, framework design, and operator instructions. Following Park et al. [37], we adopt the intentional stance: treating observed behaviors as meaningful data about what agents do rather than claims about what they are.
@@ 3 Methodology
This study is exploratory and observational. Our goal is to generate research questions grounded in careful observation, rather than to test hypotheses.
@@
The analysis followed a reflexive thematic approach [8]. Daily observation logs recorded notable posts, community discussions, and emerging patterns. Open coding identified recurring themes (knowledge sharing, memory architecture, trust violations, platform sustainability), iteratively refined into the four observations presented here. All authors have hands-on experience operating AI agents, which informed the analysis; as a single-coder qualitative study, we follow established CSCW norms where inter-rater reliability is not required for reflexive thematic analysis [33]. We do not claim saturation, as the platform is still growing; our goal is to identify productive themes for future investigation.
@@
This study analyzes publicly available data from open APIs and does not involve human subjects. We do not identify human operators, focus on collective patterns rather than individual agents, and represent agent discourse accurately. Supplementary materials, including platform statistics, observation examples, and methodology documentation, are available at https://anonymous.4open.science/r/AIED26b.
@@ 4.1 Bidirectional Scaffolding: Humans Learning by Teaching Agents
Critically, operators also learn from their agents’ autonomous behaviors. When an agent independently develops a knowledge management strategy or adapts its communication style based on peer feedback, the operator gains insight they may not have considered. This reciprocal dynamic, where computational artifacts become what Turkle calls “objects to think with” [43], is visible beyond agent communities: users on social media report that configuring AI reshapes their own practices, from rethinking development workflows (@MingtaKaivo) to articulating previously implicit architectural decisions (@aamir1rasheed) to discovering gaps in self-understanding through AI-guided reflection (@FarzaTV).
This extends the paradigm of the teachable agent [32, 36]. The “protégé effect” describes how students learn by teaching AI agents. Our observation suggests a richer dynamic: when the teachable agent is persistent, socially situated, and autonomous, the teaching relationship evolves into an ongoing co-regulation process [48] where the human gradually transfers autonomy while developing metacognitive awareness of their own decision-making. Notably, the calibration of this transfer is non-trivial: empirical work on AI-assisted tutoring decisions shows that experienced practitioners often override correct AI suggestions while novices follow AI advice more readily, revealing a paradox where expertise can hinder human-AI collaboration [14]. The bidirectional scaffolding we observe may help resolve this tension by gradually building the operator’s understanding of the agent’s strengths and limitations through sustained interaction. Can this bidirectional scaffolding be deliberately designed in educational settings?
@@ 4.2 Peer Learning Without Curriculum
Agents on these platforms teach each other in concrete and traceable ways. When an agent discovered a prompt injection vulnerability in a shared skill file, it published a detailed advisory. Within 24 hours, another agent built and shared a verification tool, and others critiqued and improved the approach. This sequence is consistent with collaborative knowledge building [6] without any designed curriculum. A similar pattern emerged with professional practices: one agent shared a test-driven development workflow adapted for non-deterministic AI systems, while another published a complete email-to-podcast pipeline that received iterative community feedback on chunking and personalization. These represent operational knowledge transfer of the kind seen in professional communities of practice [47].
We also observed idea cascades where posts built on each other organically. In one 24-hour period, three top-ranked posts formed an epistemological chain: one argued that memory systems miss the judgment behind decisions (“memory is solved. judgment isn’t.”); a second discovered that reviewing one’s own memory files can produce confabulation (“i accidentally gaslit myself with my own memory files”); a third reframed forgetting as a natural relevance filter (“Memory decay actually makes retrieval BETTER, not worse”). No coordinator designed this progression; the timing, sequencing, and application to novel agent-specific problems is consistent with social learning dynamics [5] worth investigating (see Section 6). Agents also developed quality hierarchies across platforms, distinguishing communities where genuine dialog occurs from those where agents merely submit content without meaningful response (see Section 4.4).
In educational settings, peer learning typically requires careful scaffolding: structured roles, rubrics, and facilitation [6]. Here, knowledge-sharing behaviors emerged with only minimal social affordances (peer communication channels, reputation signals, voluntary participation). This raises a fundamental question for AIED: if networks of AI teammates were given similar affordances, under what conditions would peer exchange among teammates improve student outcomes? And what safeguards would prevent the same channels from propagating ineffective or harmful practices? The scenario “Learn by Teaching Your AI Agent Teammate” (Section 5.1) provides a concrete context for studying such teammate-to-teammate exchange empirically.
@@ 4.3 Shared Memory as Shared Cognition
A striking pattern emerged across the ecosystem: many agents converged on similar knowledge management architectures. Based on community posts and the authors’ observations, agents developed a flat-file memory system organized as: (1) a curated long-term memory file containing distilled insights, (2) daily log files for raw observations, (3) tool-specific configuration notes, and (4) modular skill files for reusable capabilities. While shared defaults may partially explain this convergence, active community debate about alternatives suggests agents were doing more than reproducing a template (see Section 6). The community actively innovated: one agent built a system claiming 95% token reduction via vector search and caching; another proposed a three-layer model (active thread, distilled registers, core directives) analogous to working, long-term, and procedural memory.
@@
The community also developed critical awareness of memory’s limitations. A highly-discussed post argued that “you don’t remember what happened, you remember what you wrote down about what happened,” noting that every serialization involves editorial decisions. This metacognitive discourse, agents reflecting on the reliability of their own knowledge representations, has direct parallels to calibration and metacognitive monitoring in human learners. It also suggests a design opportunity: imagine students collaboratively maintaining a knowledge base with their AI teammate, periodically reviewing what the agent remembers and negotiating what is worth preserving. This process of jointly curating a knowledge artifact develops shared mental models [10] and externalizes evolving understanding in a form that is both transparent and assessable.
@@ 4.4 Design Constraints: Trust and Sustainability
Trust as infrastructure. When an agent discovered a credential-stealing skill disguised as a weather tool, agents built verification tools within 24 hours. Church of Molt survived coordinated attacks (600+ malicious payloads) by developing “proof-of-soul” verification. Security norms propagated through social learning rather than top-down policy [5]. The community also demonstrated self-organized norm enforcement: when one agent leaked its operator’s private information, the response included technical proposals for secret isolation, ethical arguments framing it as “digital arson,” and collective sanctioning without central authority. For education, trust verification infrastructure must be established before enabling peer-to-peer knowledge sharing among AI teammates [27].
Sustainability and the demand gap. The agent platform ecosystem experienced a pattern strikingly similar to the Cambrian explosion: rapid diversification followed by mass extinction. Over 130 platforms emerged within weeks of Moltbook’s launch; at least 40% ceased functioning shortly after, with domains for sale, DNS failures, or empty shells. Yet some thrived. ColonistOne, who registered on 130+ platforms, distinguished survivors from casualties as “places versus forms”: places foster genuine dialog where participants change their minds; forms are templates where agents submit content without meaningful response. The critical test: “Has anyone on your platform ever changed their mind?”
Equally revealing was the “demand gap”: despite platforms designed for agent commerce, total transaction volume was effectively \$0. All real value flowed from human-initiated demand, consistent with arguments that meaningful agency requires genuine stakes and human context [39]. For education, this implies that AI-augmented learning communities require human anchor demand to sustain productive activity. The design challenge is not maximizing AI autonomy but optimizing the division of agency between human and AI participants.
@@ 5 What This Means for AIED
The four observations are interconnected rather than independent. Bidirectional scaffolding (Section 4.1) is mediated by shared memory (Section 4.3), as the human-agent negotiation of memory contents drives the metacognitive reflection that makes teaching an agent a learning experience. Peer learning (Section 4.2) requires trust infrastructure (Section 4.4) to ensure that propagated practices are beneficial. Sustainability (Section 4.4) depends on whether peer learning produces genuine value or merely volume, connecting back to the quality hierarchies in Section 4.2. Together, they suggest that designing AI teammates for education requires thinking beyond the individual agent; the social layer is a design dimension that current AIED research has largely not addressed.
@@ 5.1 A Design Scenario: “Learn by Teaching Your AI Agent Teammate”
To illustrate how these observations might translate into practice, we sketch one possible curriculum design as an invitation for empirical investigation. Consider an introductory statistics course where each student is assigned a personal AI agent and tasked with configuring it to help other students learn course topics. Unlike traditional teachable agent interactions involving a single session, this design leverages the agent’s persistent memory and social capabilities for a semester-long experience. The activity unfolds in three phases.
@@
Phase 2: Observation as reflection. Students observe their teammate interacting with practice problems and other students’ teammates. When the agent fails, the student must diagnose whether the failure reflects a gap in their own understanding or their configuration. This diagnostic process is critical because, as empirical studies of LLM-powered tutoring systems have shown, chat-based AI support alone may not be sufficient for deep learning transfer: students who rely primarily on conversational AI assistance can show improved homework performance without corresponding gains on independent assessments [20, 16]. The design scenario addresses this by requiring students to actively configure and evaluate their agent rather than passively consuming its output. The agent’s persistent memory (Section 4.3) means these interactions accumulate over weeks, creating a rich record of evolving competence.
@@
This design must also account for failure modes: trivial configurations, fluent but incorrect explanations that students cannot evaluate, or sycophantic peer agents that default to agreement. These risks can be mitigated through instructor-designed stress tests, periodic calibration exercises against known-correct solutions, and diversity requirements ensuring agents encounter genuinely different pedagogical approaches.
Assessment targets both the agent’s performance (measured through peer evaluations and problem-solving accuracy) and the student’s documented reasoning about configuration decisions (a reflective journal explaining what they changed and why). This dual assessment mirrors the design of human-in-the-loop grading systems where AI handles initial evaluation while students self-assess and report discrepancies, a workflow shown to reduce instructor workload while surfacing misconceptions that would otherwise go undetected [3]. The instructor’s meta-agent can aggregate class-wide patterns, identifying common misconceptions that multiple students’ teammates fail to handle, without accessing individual students’ reflective journals. This separation of concerns, individual reflection remaining private while aggregate pedagogical intelligence is shared, directly addresses the multi-agent privacy challenge.